Factor
1. a^2+3a+8a+24
2. 3u^2+1u-24
3. 28v^2-37v-21
4. 9g^2+18gp+8p^2

Answer by stanbon(57342)

Factor
1. a^2+3a+8a+24
---
= a(a+3) + 8(a+3)
= (a+3)(a+8)
---------------------
2. 3u^2+1u-24
= 3u^2+9u-8u-24
------
= 3u(u+3)-8(u+3)
---
= (u+3)(3u-8)
----------------------
3. 28v^2-37v-21
a = 28 ; b = -37 ; c = -21
No simple factors.
----------
4. 9g^2+18gp+8p^2
= (3q+2p)(3q+4p)
===================
Cheers,
Stan H.
===================